OpenStudy (anonymous):

Given f(x)= 1/3x^2 and g(x)=square root x+4. Find f of g(x)

OpenStudy (anonymous):

by f of g(x) you mean f(g(x)...

OpenStudy (anonymous):

nwo g(x) = sqrt(x+4) bcoz f(x) = 1/3x^2 so f(sqrt(x+4)) = 1/3(sqrt(x+4))^2 = 1/3(x+4)

OpenStudy (anonymous):